OpenStudy (anonymous):

Simplify. -14x^3/x^3-5x^4 from connections acadamy i sck at this

OpenStudy (anonymous):

\[\frac{ -14x^3}{x^3-5x^4} \] like that?

OpenStudy (anonymous):

factor out a \(x^3\) from the denominator, then cancel it

OpenStudy (anonymous):

i still dont get how to do it im rlly ad at math

OpenStudy (anonymous):

\[\frac{ -14x^3}{x^3-5x^4}=\frac{-14x^3}{x^3(1-5x)}=\frac{-14}{1-5x}\]

OpenStudy (anonymous):

-14\[-14/{1-5x}\]

OpenStudy (anonymous):

so the answer would be: -14/1-5x;where x is not equal to 1/5